# Effective Interplay between Sparsity and Quantization: From Theory to Practice

### v1

@@ 4 Experimental methodology and results
Models, datasets, and evaluation setup. We study the most widely adopted Transformer-based models, including OPT [69] and LLaMA [58] model families. In line with prior work [64, 15, 53], we fine-tune pretrained models and evaluate perplexity on the WikiText2 [36] dataset. In addition, we assess orthogonality across different metrics of ViT [10] on ImageNet-1k [7]. In all experiments, we designate the dense FP32 configuration as our primary baseline. Our experiments span a diverse range of configurations to validate our theoretical findings, including various variants of max-scaled formats, such as INT8 quantization with per channel scaling [8], HBFP8/6 [11], and MXFP8/6 [50]. We primarily study magnitude-based 50% unstructured and 2:4 structured sparsity with sparsity-aware fine-tuning. We evaluate the impact of increased compression on ViT-B/16 by applying 75% unstructured sparsity, 1:4 structured sparsity, and HBFP4. Detailed results can be found in Appendix F. We also explore post-training one-shot sparsity methods like SparseGPT [15] and Wanda [53] in Appendix E. In our experiments, we evaluate the impact of each compression method by analyzing the variations in perplexity and cross-entropy loss compared to the baseline, thereby focusing on the cumulative error in the final model output. Additionally, we analyze the errors of intermediate layers to support our mathematical analysis, as detailed in Appendix G. Note that we focus on cross-entropy loss for ViT-B/16 instead of classification accuracy. This is because accuracy remains unaffected as long as the most likely label remains unchanged, regardless of its absolute value. However, our primary metric for orthogonality bound is the aggregated errors introduced in the final model output distribution. Table 7 in Appendix F evaluates orthogonality bound on additional metrics.
@@
Experimental setup. We exclusively sparsify and/or quantize layers with trainable parameters. Specifically, we target all linear layers in LLMs, and all linear and convolution layers (including the initial embedding layer) in ViT-B/16. These layers collectively constitute approximately 99% of the total parameters. In all experiments, we sparsify weights while keeping activation dense. Both weights and activation tensors are quantized before matrix multiplication operations. For OPT, LLaMA, and ViT fine-tuning, we employ sparse fine-tuning on a dense FP32 pre-trained model, recomputing sparsity masks at each iteration. In experiments involving sparsity followed by quantization (SQ), we apply one-shot quantization to sparse fine-tuned models. Conversely, for experiments with the reverse order (QS), we directly fine-tune the model in a quantized and sparsified manner. At each iteration, we quantized activations and weights while applying sparsity to weight tensors. We validate the effectiveness of these compression recipes through an ablation study, the details of which are presented in Appendix D. To ensure fair comparison, we maintain uniform hyperparameters across various number formats for a given model and sparsity type (details in Appendix B). We present a limited sensitivity study on the initial seed number in Table 8 in Appendix H.
@@ 4.1 Empirical study 1: Order of sparsity and quantization
This section presents empirical evidence demonstrating that applying sparsity before quantization leads to better perplexities compared to the reverse order. These results align with the theoretical arguments in Section 3.1. Table 1 presents perplexities for OPT-125M and LLaMA-2-7B under various number formats and sparsity types. We compare the order of operations across different configurations: sparsity followed by quantization (SQ) and quantization followed by sparsity (QS). Configurations with lower perplexities are highlighted in bold. Consistently, we observe that the SQ order yields better perplexities across all number formats for magnitude-based sparsity. As discussed in Section 3.1, in the case of QS, quantizing a tensor can alter the order of its elements due to changes in magnitudes. If magnitude-based sparsity removes elements of the quantized tensor that were originally larger before quantization, the error of the composition may exceed the sum of the errors caused by each transformation individually. This compounded error further propagates through subsequent dot-product and vector operations, impacting overall model performance.
@@ 4.2 Empirical study 2: Non-Orthogonality between sparsity and quantization
Sparsity type | Number format | PPL | Orth. bound | PPL | Orth. bound | PPL | Orth. bound | PPL | Orth. bound | CE Loss | Orth. bound |
@@
50% | FP32 | 29.94 | - | 11.30 | - | 6.31 | - | 10.09 | - | 0.723 | - |
INT8 | 30.22 | 30.35 | 11.37 | 11.39 | 6.94 | 6.38 | 10.85 | 10.19 | 0.728 | 0.725 | |
MXFP8 | 31.13 | 30.74 | 11.74 | 11.69 | 6.40 | 6.36 | 10.34 | 10.18 | 0.745 | 0.742 | |
MXFP6 | 31.20 | 30.30 | 11.53 | 11.44 | 6.38 | 6.35 | 10.15 | 10.18 | 0.734 | 0.735 | |
HBFP8 | 30.46 | 30.18 | 11.31 | 11.32 | 6.32 | 6.31 | 10.12 | 10.12 | 0.724 | 0.723 | |
HBFP6 | 32.51 | 32.2 | 11.94 | 11.65 | 6.51 | 6.43 | 10.55 | 10.43 | 0.736 | 0.737 | |
2:4 | FP32 | 31.89 | - | 15.48 | - | 9.30 | - | 13.07 | - | 0.759 | - |
INT8 | 32.76 | 32.30 | 15.61 | 15.57 | 9.37 | 9.37 | 13.23 | 13.17 | 0.762 | 0.761 | |
MXFP8 | 33.99 | 32.69 | 15.70 | 15.87 | 9.35 | 9.35 | 13.35 | 13.16 | 0.781 | 0.777 | |
MXFP6 | 33.41 | 32.25 | 15.95 | 15.64 | 9.32 | 9.34 | 13.2 | 13.16 | 0.770 | 0.771 | |
HBFP8 | 32.25 | 32.05 | 15.57 | 15.50 | 9.39 | 9.31 | 13.11 | 13.1 | 0.760 | 0.759 | |
HBFP6 | 34.58 | 34.15 | 16.98 | 15.82 | 10.68 | 9.42 | 13.64 | 13.41 | 0.774 | 0.773 |
This section demonstrates that combining sparsity and quantization results in additional error, surpassing the sum of their individual errors. Table 2 presents perplexities for OPT-125M, OPT-6.7B, LLaMA-2-7B and LLaMA-3-8B on WikiText2, and cross-entropy (CE) loss for ViT-B/16 on ImageNet-1k for various combinations of number formats and sparsity schemes. Following our conclusions regarding the order of operations, we only report results for SQ. We compute the orthogonality bound for each combination by summing the individual errors from quantization and sparsity relative to the baseline dense FP32 model, using Equation 26. Each model’s performance is compared against this bound, with superior results highlighted in bold. In the majority of configurations, perplexity and cross-entropy loss values exceed the orthogonality bounds, validating the non-orthogonality of sparsity and quantization.
@@
We mathematically show (Section 3.2) that the additional error introduced by combining sparsity and quantization significantly depends on the values of quantized activation tensors and the quantization error of sparsified weights. This error tends to amplify through successive dot products and vector operations. Consequently, the quantization error affects the additional error caused by the composition more than the sparsification error. The gap between the model’s performance and the orthogonality bound is minimal for number formats with minimal performance decrease, whereas larger errors are observed for formats with larger errors. For instance, HBFP6 results in a 2.26 increase in perplexity for OPT-125M, while its combination with 50% unstructured sparsity leads to a 4.86 increase.
@@
Our analysis reveals that both model size and compression ratio significantly influence the additional error introduced by combining sparsity and quantization. Larger models exhibit greater tolerance to these compression methods, resulting in lower additional errors. Moreover, formats with minimal quantization errors (e.g. MXFP8), and sparsity types with minimal sparsification errors (e.g. unstructured sparsity) lead to lower additional errors, even for smaller models. The effect of high quantization error is more pronounced for sparsity schemes known for higher errors. For instance, HBFP6’s combination with 2:4 sparsity causes a 6.93 perplexity increase for OPT-125M and a 6.12 perplexity increase for OPT-6.7B. In contrast, combining HBFP6 with unstructured sparsity results in smaller increases of 4.86 and 0.79 for the respective models.
@@
We also observe instances where the orthogonality bound is slightly higher than the actual perplexity: (a) INT8 with 50% unstructured sparsity for both OPT models, (b) MXFP8 with 2:4 sparsity for OPT-6.7B, (c) MXFP6 with 2:4 sparsity for LLaMA-2-7B, and (d) MXFP6 50% unstructured sparsity for LLaMa-3-8B. Although these occurrences do not consistently correlate with specific formats, sparsity types, or model sizes, they do not contradict our theoretical analysis, which primarily concerns upper bounds of errors and does not entirely rule out orthogonal cases. Furthermore, orthogonal configurations still result in larger errors compared to applying either sparsity or quantization alone. Our mathematical analysis (Theorem 3.9) indicates that there exists at least one occurrence where the orthogonality is not preserved. This underscores the need for careful examination when applying these compression methods together, as they do not guarantee high accuracies. Delving into cases where orthogonality is preserved falls beyond the scope of this paper and we leave it for future work.
@@
We also observe that although the cross-entropy loss for ViT-B/16 is higher than the calculated orthogonality bound in most cases in Table 2, the difference is relatively small. We hypothesize the reason behind this behavior is due to the fact that ViT-B/16 being a vision model that operates on images is more robust to quantization and sparsity errors than LLMs that operate on text. Hence, the quantization and sparsity levels shown in Table 2 are not sufficient to induce large errors, and hide the non-orthogonality of quantization and sparsity. To test our hypothesis, we increase the compression further by employing 75% unstructured sparsity, 1:4 structured sparsity and HBFP4 on ViT-B/16 (Appendix F). The results show that the difference between the baseline cross-entropy loss and the calculated orthogonality bound for these cases increases significantly validating the non-orthogonality of quantization and sparsity. This increase in cross-entropy loss translates to a significant non-orthogonal drop in accuracy as well.
@@ 5 Discussion
Our theoretical analysis and experimental results offer multiple insights for ML model practitioners. First, our analysis demonstrates a risk-free method to improve model performance, measured by lower perplexity and/or higher accuracy, through choosing the optimal ordering of compression operations for any max-scaled number format and magnitude-based pruning scheme. This contribution is particularly important in the current ML landscape, where sparsity and quantization are pivotal methods for reducing the memory footprint and bandwidth requirements of state-of-the-art LLMs. Second, we show that calculating the orthogonality bound offers a close enough estimate of model performance (e.g., accuracy, perplexity, etc.) under conditions of quantization and sparsity. This bound can streamline the search for optimal sparse-quantized model configurations by effectively narrowing the search space. There is an inherent tradeoff between the hardware benefits of various sparse-quantized configurations and the achieved model performance. Quantization bit-width and sparsity level are key factors influencing the memory and bandwidth requirements for serving these models. For example, at a 50% sparsity level, 8-bit and 6-bit quantization result in total reductions in memory footprint and bandwidth requirements by 8 and 10.7, respectively.
Ideally, practitioners aim to maximize compression (increase sparsity ratio and/or reduce quantization effective bitwidth. Our analysis elucidates the individual and combined impacts of these factors across a range of recent large models, providing practical guidelines to achieve the highest compression without compromising model performance. Typically, 8-bit quantization with any max-scaled number format can serve as a direct replacement for FP32 when combined with any form of sparsity in the optimal order (SQ). As discussed in Section 4, certain models exhibit sensitivity to sub-8bit number formats and structured sparsity combinations, even when applied in an optimal order. In scenarios where improvements in arithmetic density (TOPS/mm2) and memory footprint justify a slight reduction in model performance, such as the deployment of large models on edge devices, these combinations may still be viable.
In this work, we do not consider heterogeneous sparsity and quantization schemes, where the sparsity fraction and quantization bit-width vary across layers and differ between activation and weight tensors. Such approaches [50, 35, 20, 12] have demonstrated to be effective in maintaining model accuracy or perplexity while improving compressing ratio. However, these schemes may not be hardware-friendly, introduce noticeable overhead, and are impractical to implement on off-the-shelf hardware platforms (e.g., GPU, TPU). We leave the investigation of the interactions between these heterogeneous quantization and sparsity schemes for future work.

### v2

@@ 4 Experimental methodology and results
Models, datasets, and evaluation setup. We study the most widely adopted Transformer-based models, including OPT (Zhang et al., 2022b) and LLaMA (Touvron et al., 2023) model families. In line with prior work (Xiao et al., 2023; Frantar & Alistarh, 2023; Sun et al., 2023), we fine-tune pre-trained models and evaluate perplexity on the WikiText2 (Merity et al., 2017) dataset. The pretrained LLMs used in our experiments are base (general-purpose) models, not instruct-tuned variants. In addition, we assess non-orthogonality across different metrics of ViT (Dosovitskiy et al., 2021) and ResNet (He et al., 2016) on ImageNet-1k (Deng et al., 2009). In all experiments, we designate the dense FP32 configuration as the primary baseline.
Our experiments span a diverse range of configurations to validate our mathematical findings, including various variants of max-scaled formats, such as INT8 quantization with per channel scaling (Dettmers et al., 2022), HBFP8/6 (Drumond et al., 2018a), and MXFP8/6 (Rouhani et al., 2023b). We primarily study magnitude-based 50% unstructured and 2:4 structured sparsity with sparsity-aware fine-tuning. We define N:M structured sparsity as following: in every group of M consecutive weights, at most N weights can have non-zero values. We evaluate the impact of a higher compression ratio on ViT-B/16 and ResNet-50 by applying 75% unstructured sparsity, 1:4 structured sparsity, and HBFP4. Detailed results can be found in Appendix G and O. We also explore post-training one-shot sparsity methods like SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2023) in Appendix F.
In our experiments, we evaluate the impact of each compression method by analyzing the variations in perplexity and cross-entropy loss compared to the baseline, thereby focusing on the cumulative error in the model output. Additionally, we analyze the errors of intermediate layers to support our mathematical analysis, as detailed in Section 4.3 and Appendix I.
Note that we focus on cross-entropy loss for ViT-B/16 and ResNet-50 instead of classification accuracy. This is because accuracy remains unaffected as long as the most likely label remains unchanged, regardless of its absolute value. However, our primary metric for orthogonality threshold is the aggregated errors introduced in the model output distribution. Table 8 in Appendix G and Table 12 in Appendix O present orthogonality threshold on additional metrics.
@@
Experimental setup. We exclusively sparsify and/or quantize layers with trainable parameters. Specifically, we target all linear layers in LLMs (excluding the lm-head or embedding layers following the literature (Frantar et al., 2022; Frantar & Alistarh, 2023; Lee et al., 2024)), and all linear and convolution layers (including the initial embedding layer) in ViT-B/16 and ResNet-50. These layers collectively constitute approximately 99% of the total parameters. In all experiments, we sparsify weights while keeping activation dense. Both weights and activation tensors are quantized before matrix multiplication operations. For OPT, LLaMA, ViT, and ResNet fine-tuning, we employ sparse fine-tuning on a dense FP32 pre-trained model, recomputing sparsity masks at each iteration. In experiments involving sparsity followed by quantization (SQ), we apply one-shot quantization to sparse fine-tuned models. Conversely, for experiments with the reverse order (QS), we directly fine-tune the model in a quantized and sparsified manner. At each iteration, we quantize activations and weights while applying sparsity to weight tensors. We validate the effectiveness of these compression recipes through an ablation study, the details of which are presented in Appendix D. To ensure fair comparison, we maintain uniform hyperparameters across various number formats for a given model and sparsity type (details in Appendix B). We present a limited sensitivity study on the initial seed number in Table 9 in Appendix H. A summary of our experimental setup is presented in Table 6.
@@ 4.1 Empirical study 1: Order of sparsity and quantization
This section presents empirical evidence demonstrating that applying sparsity before quantization leads to better perplexities compared to the reverse order. These results are aligned with the mathematical analysis in Section 3.1. Table 1 presents perplexities for OPT-125M and LLaMA-2-7B under various number formats and sparsity types considering both orders of transformations. “50%” denotes unstructured sparsity, while “2:4” represents a variant of N:M structured sparsity. In the FP32 columns, only one perplexity for each compression order is reported because no quantization is applied. The best results for each (sparsity type, number format) pair are highlighted in bold.
We evaluate the order of transformations across different configurations: sparsity followed by quantization (SQ) and quantization followed by sparsity (QS). Configurations with lower perplexities are highlighted in bold. Consistently, we observe that the SQ order yields better perplexities across all number formats for magnitude-based sparsity. As discussed in Section 3.1, in the case of QS, quantizing a tensor can alter the order of its elements due to changes in magnitudes. If magnitude-based sparsity removes elements of the quantized tensor that were originally larger before quantization, the error of the combination may exceed the sum of the errors caused by each transformation individually. This compounded error further propagates through subsequent dot-product and vector operations, impacting overall model performance.
This section demonstrates that combining sparsity and quantization results in additional error, surpassing the sum of their individual errors. Table 2 presents perplexities for OPT-125M, OPT-6.7B, LLaMA-2-7B and LLaMA-3-8B on WikiText2, and cross-entropy (CE) loss for ViT-B/16 on ImageNet-1k for various combinations of number formats and sparsity types. Following our conclusions regarding the order of transformations, we only report results for SQ. We compute the orthogonality threshold for each combination by summing the individual errors from sparsity and quantization relative to the baseline dense FP32 model, using Equation 12. Each model’s performance is compared against this bound, with superior results highlighted in bold. In the majority of configurations, perplexity and cross-entropy loss values exceed the orthogonality thresholds, validating the non-orthogonality of sparsity and quantization.
@@ 4.2 Empirical study 2: Non-Orthogonality between sparsity and quantization
Sparsity type | Number format | PPL | Orth. threshold | PPL | Orth. threshold | PPL | Orth. threshold | PPL | Orth. threshold | CE Loss | Orth. threshold |
@@
50% | FP32 | 29.94 (+2.29) | - | 11.30 (+0.44) | - | 6.31 (+1.19) | - | 10.09 (+4.56) | - | 0.723 (+0.020) | - |
INT8 | 30.22 (+2.16) | 30.35 | 11.37 (+0.42) | 11.39 | 6.94 (+1.75) | 6.38 | 10.85 (+5.22) | 10.19 | 0.728 (+0.022) | 0.725 | |
MXFP8 | 31.13 (+2.68) | 30.74 | 11.74 (+0.49) | 11.69 | 6.40 (+1.23) | 6.36 | 10.34 (+4.72) | 10.18 | 0.745 (+0.023) | 0.742 | |
MXFP6 | 31.20 (+3.19) | 30.30 | 11.53 (+0.51) | 11.44 | 6.38 (+1.22) | 6.35 | 10.15 (+4.53) | 10.18 | 0.734 (+0.019) | 0.735 | |
HBFP8 | 30.46 (+2.65) | 30.18 | 11.31 (+0.43) | 11.32 | 6.32 (+1.20) | 6.31 | 10.12 (+4.56) | 10.12 | 0.724 (+0.020) | 0.723 | |
HBFP6 | 32.51 (+2.60) | 32.2 | 11.94 (+0.74) | 11.65 | 6.51 (+1.27) | 6.43 | 10.55 (+4.68) | 10.43 | 0.736 (+0.018) | 0.737 | |
2:4 | FP32 | 31.89 (+4.24) | - | 15.48 (+4.62) | - | 9.30 (+4.18) | - | 13.07 (+7.54) | - | 0.759 (+0.056) | - |
INT8 | 32.76 (+4.70) | 32.30 | 15.61 (+4.66) | 15.57 | 9.37 (+4.18) | 9.37 | 13.23 (+7.60) | 13.17 | 0.762 (+0.056) | 0.761 | |
MXFP8 | 33.99 (+5.54) | 32.69 | 15.70 (+4.45) | 15.87 | 9.35 (+4.18) | 9.35 | 13.35 (+7.73) | 13.16 | 0.781 (+0.059) | 0.777 | |
MXFP6 | 33.41 (+5.40) | 32.25 | 15.95 (+4.93) | 15.64 | 9.32 (+4.16) | 9.34 | 13.20 (+7.58) | 13.16 | 0.770 (+0.055) | 0.771 | |
HBFP8 | 32.25 (+4.44) | 32.05 | 15.57 (+4.69) | 15.50 | 9.39 (+4.27) | 9.31 | 13.11 (+7.55) | 13.1 | 0.760 (+0.056) | 0.759 | |
HBFP6 | 34.58 (+4.67) | 34.15 | 16.98 (+5.78) | 15.82 | 10.68 (+5.44) | 9.42 | 13.64 (+7.77) | 13.41 | 0.774 (+0.056) | 0.773 |
@@
We mathematically show (Section 3.2) that the additional error introduced by combining sparsity and quantization significantly depends on the values of quantized activation tensors and the quantization error of sparsified weights. This error tends to amplify through successive dot products and vector operations. Consequently, the quantization error affects the additional error caused by the combination more than the sparsity error. The gap between the model’s performance and the orthogonality threshold is minimal for number formats with minimal performance decrease, whereas larger errors are observed for formats with larger errors. For instance, HBFP6 results in a 2.26 increase in perplexity for OPT-125M, while its combination with 50% unstructured sparsity leads to a 4.86 increase.
@@
Our analysis reveals that both model size and compression ratio significantly influence the additional error introduced by combining sparsity and quantization. Larger models exhibit greater tolerance to these compression methods, resulting in lower additional errors. Moreover, formats with minimal quantization errors (e.g. MXFP8), and sparsity types with minimal sparsification errors (e.g. unstructured sparsity) lead to lower additional errors, even for smaller models. The effect of high quantization error is more pronounced for sparsity types known for higher errors. For instance, HBFP6’s combination with 2:4 sparsity causes a 6.93 perplexity increase for OPT-125M and a 6.12 perplexity increase for OPT-6.7B. In contrast, combining HBFP6 with unstructured sparsity results in smaller increases of 4.86 and 0.79 for the respective models.
@@
We also observe instances where the orthogonality threshold is slightly higher than the actual perplexity: (a) INT8 with 50% unstructured sparsity for both OPT models, (b) MXFP8 with 2:4 sparsity for OPT-6.7B, (c) MXFP6 with 2:4 sparsity for LLaMA-2-7B, and (d) MXFP6 50% unstructured sparsity for LLaMa-3-8B. Although these occurrences do not consistently correlate with specific formats, sparsity types, or model sizes, they do not contradict our mathematical analysis, which primarily concerns upper bounds of errors and does not entirely rule out orthogonal cases. Furthermore, orthogonal configurations still result in larger errors compared to applying either sparsity or quantization alone. Our mathematical analysis (Theorem 3.10) indicates that there exists at least one occurrence where the orthogonality is not preserved. This underscores the need for careful examination when applying these compression methods together, as they do not guarantee high accuracies. Delving into cases where orthogonality is preserved falls beyond the scope of this paper, and we leave it for future work.
@@
We observe that although the cross-entropy loss for ViT-B/16 is higher than the calculated orthogonality threshold in most cases in Table 2, the difference is relatively small. We hypothesize the reason behind this behavior is due to the fact that ViT-B/16, being a vision model that operates on images, is more robust to sparsity and quantization errors than LLMs that operate on text. Hence, the sparsity and quantization levels shown in Table 2 are not sufficient to induce large errors, and hide the non-orthogonality of sparsity and quantization. To test our hypothesis, we increase the compression further by employing 75% unstructured sparsity, 1:4 structured sparsity and HBFP4 on ViT-B/16 (Appendix G). The results show that the difference between the baseline cross-entropy loss and the calculated orthogonality threshold for these cases increases significantly, validating the non-orthogonality of sparsity and quantization. This increase in cross-entropy loss also translates to a significant non-orthogonal drop in accuracy. Additionally, we run the same experiments on a CNN-based vision model, ResNet-50 (Appendix O), and the experimental results corroborate these findings.
4.3 Ablation: Error Propagation Across Layers
In this section, we study error propagation across the layers of a deep neural network.
While our mathematical study provides insights into how sparsity and quantization introduce errors at the dot-product and tensor levels, mathematically estimating the cumulative effect of these errors on the model’s final loss or performance is non-trivial. Therefore, we perform this layer-wise empirical analysis to verify that the order of applying sparsity and quantization not only affects per-layer errors but also significantly impacts overall model’s accuracy. By inspecting the intermediate outputs of the pre-trained OPT-125M model, we show that the error accumulation increases with the layer index and that the choice of the order directly influences the final model performance.
We apply quantization and magnitude-based sparsity to all linear layers of the pre-trained OPT-125M model in a zero-shot manner. We feed a sample from the test subset through both the compressed model and the corresponding full precision dense model, and we measure the difference in Feed-Forward outputs for each Transformer block. We repeat this experiment for both the SQ and QS orders, and we compare the errors at each layer. The results are shown in Figure 1.
As illustrated in the figure, the per-layer error increases consistently with the layer index, regardless of the order of transformations, and reaches its peak in the final layer. However, we observe that the SQ order produces significantly lower errors at each intermediate layer compared to the reverse order. This pattern indicates that the error introduced by QS accumulates more rapidly as layers deepen, highlighting the detrimental effects of applying quantization before sparsity.
@@ 5 Discussion
Our mathematical analysis and experimental results offer multiple insights for ML model practitioners. First, our analysis demonstrates a risk-free method to improve model performance, measured by lower perplexity and/or higher accuracy, through choosing the optimal ordering of compression operations for any max-scaled number format and magnitude-based pruning scheme. This contribution is particularly important in the current ML landscape, where sparsity and quantization are pivotal methods for reducing the memory footprint and bandwidth requirements of state-of-the-art LLMs. Second, we show that calculating the orthogonality threshold offers a close enough estimate of model performance (e.g. accuracy, perplexity, etc.) under conditions of sparsity and quantization. This bound can streamline the search for optimal sparse-quantized model configurations by effectively narrowing the search space. There is an inherent tradeoff between the hardware benefits of various sparse-quantized configurations and the achieved model performance. Quantization bit-width and sparsity level are key factors influencing the memory and bandwidth requirements for serving these models. For example, at a 50% sparsity level, 8-bit and 6-bit quantization result in total reductions in memory footprint and bandwidth requirements by 8 and 10.7, respectively.
Ideally, practitioners aim to maximize compression (increase sparsity ratio and/or reduce the average bitwidth per element). Our analysis elucidates the individual and combined impacts of these factors across a range of recent large models, providing practical guidelines to achieve the highest compression without compromising model performance. Typically, 8-bit quantization with any max-scaled number format can serve as a direct replacement for FP32 when combined with any form of sparsity in the optimal order (SQ). As discussed in Section 4, certain models exhibit sensitivity to sub-8bit number formats and structured sparsity combinations, even when applied in an optimal order. In scenarios where improvements in arithmetic density (TOPS/mm2) and memory footprint justify a slight reduction in model performance, such as the deployment of large models on edge devices, these combinations may still be viable.
In this work, we do not consider heterogeneous sparsity and quantization schemes, where the sparsity fraction and quantization bit-width vary across layers and differ between activation and weight tensors. Such approaches (Rouhani et al., 2023b; Ma et al., 2024; Harma et al., 2022; Evci et al., 2020b) have demonstrated to be effective in maintaining model accuracy or perplexity while improving compressing ratio. However, these schemes may not be hardware-friendly, introduce noticeable overhead, and are impractical to implement on off-the-shelf hardware platforms (e.g. GPU, TPU). We leave the investigation of the interactions between these heterogeneous sparsity and quantization schemes for future work.